OpenStudy (anonymous):

evaluate the integral (x^2 + 3)e^(3x+5) dx

OpenStudy (anonymous):

Are you supposed to use U substitution?

OpenStudy (anonymous):

x^2e^(3x+5) + 3e^(3x+5) for the 3e^(3x+5) sub 3x + 5 = u then du = 3dx and i guess you can do it for x^2e^(3x+5) you should do integration by parts

OpenStudy (anonymous):

So would the answer be (x+3) e(3x+5) ???

OpenStudy (anonymous):

I meant e^(3x+5)

OpenStudy (anonymous):

U= x^2 dV = e^(3x+5) dU =2x V= 1/3 e^(3x+5) \[\int\limits_{ }^{ } x ^{2}e ^{3x+5} dx = x ^{2}\frac{ 1 }{ 3 }e ^{3x+5} - \int\limits_{ }^{ } 2x\frac{ 1 }{ 3 }e ^{3x+5}dx\]

OpenStudy (anonymous):

see how the integral on the RHS is simpler than the one on the LHS?

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

Yes, I see